OpenStudy (anonymous):

An equation of the circle whose center is (-3,1) and whose radius is 8 is 1)(x-3)2+(y+1)2=64 2)(x-3)2+(y+1)2=8 3)(x+3)2+(y-1)2=64 2)(x+3)2+(y-1)2=8

OpenStudy (anonymous):

)(x-3)2+(y+2)2=4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

It should be 2)(x-3)2+(y+2)2=4

OpenStudy (reemii):

If the center is \((a,b)\), It's \((x-a)^2 + (y-b)^2= \text{radius}^2=8^2\). So the answer is \((x+3)^2+(y-1)^2=64\).

OpenStudy (anonymous):

thank you